A BUILDER IS ADDING A20'X28' ROOM TO AN EXISTING HOUSE. THE FLOORING CHOICES ARE AN 8"X8" TILE THAT COST $0.98 PER TITLE ,OR A 12"X12" TITLE THAT COST $1.98 PER TITLE. APPROX HOW MUCH WILL THE BUILDER SAVE BY USING THE LESS EXPENSIVE OPTION? NEED TO KNOW HOW TO SOLVE THIS PROBLEM
Answer by checkley75(3666)

20X28=560 SQUARE FEET TO BE COVERED.
AN 8x8 IN TILE=64 INCHES SQUARE OR 64/144=4/9 SQUARE FOOT.
A 12x12 TILE=1 SQUARE FOOT
[560/(4/9)]*.98=(560*9/4)*.98=1260*.98=$1234.80 COST USING 8X8 TILES
(560/1)*1.98=$1108.80 COST USING THE 12X12 TILES.
COST SAVING BY USING THE LARGER TILES IS:
1234.80-1108.80=&126.00 IS THE SAVINGS.